### v1

# Are EU low-carbon structural funds efficient in reducing emissions?

Abstract
This paper investigates the effectiveness of the “low-carbon economy” expenditures from European Structural and Investment Funds in fostering reductions in greenhouse gas emissions within European regions, focusing on the 2007-2013 and 2014-2020 programme periods. By decomposing emissions time series into trend and cycle components and considering them within a panel data framework, our research highlights that the impacts of low-carbon economy expenditures vary, qualitatively and quantitatively, with the targeted regions’ development level. We find significant emissions reductions in developed and transition regions yet less favourable outcomes in less developed areas. Further analysis into specific greenhouse gas emissions types (CO2, CH4, and N2O) reveals inconsistent impacts, underscoring the complexity of achieving emissions reductions. Our findings emphasise the need for tailored environmental strategies that accommodate the economic disparities of regions in the European Union.
Keywords: Greenhouse gas emissions; Emission decomposition; Low-carbon economy; European Structural and Investment Funds; Regional development
1 Introduction
The European Union (EU) has the objective of achieving climate-neutrality by 2050. This implies large-scale investments in the energy transition (McCollum et al., 2018; Klaassen and Steffen, 2023). According to the European Commission, energy investments in the EU will have to reach €396 billion per year from 2021 to 2030 and €520-575 billion per year in the subsequent decades until 2050 (European Parliament, 2023). A substantial share of these ought to be public investments that foster the development of infrastructure, the deployment of technological innovations, and the scale-up of private capital (Carraro et al., 2012; Owen et al., 2018; Meckling et al., 2022). Accordingly, the EU has enacted climate spending targets amounting to 20 of its 2014-2020 budget and 30 of its 2021-2027 budget that ought to be devoted to climate action111See European Court of Auditors (2022), which also emphasises that climate spending over the 2014-2020 period has been over-estimated.. In this context, one can ask a necessary question about the efficiency of European climate policy: have these expenses effectively led to the reduction of greenhouse gas (GHG) emissions?
This paper exploits regional variations in European climate spending to address this issue. The main investment vehicles of the EU are the European Structural and Investment Funds (ESI funds), which accounted for €535 billion over the period 2014-2020 (European Commission, 2023). By construction, these funds are spent heterogeneously. EU regions are grouped into three categories based on their relative development level (least developed regions, transition regions, and developed regions), and this determines their eligibility and their financing rate for structural funds. From the 2014-2020 programming period onwards, the usage of these funds has been tagged with a “thematic objective”. In particular, the “Low-Carbon Economy” tag allows the identification of the funds that have been assigned to climate action. Although the programming period 2007-2013 did not define thematic objectives in the same way, the aims of the expenditures are also tagged by “priority codes”, from which it is possible to determine the funds related to climate action consistently with the low-carbon economy thematic objective from 2014-2020.
Consequently, we use the total amount spent on the low-carbon economy objective in the last two budgeting periods (2007-2013 and 2014-2020) as a measure of public climate investment. We then employ an econometric panel data model to investigate the influence of these investments on GHG emissions in European regions between 2007 and 2022 at the NUTS 2 level. Our methodology involves decomposing the emission time series of each region into trend and cycle components. This strategy gives us deeper insights into the nexus between economic growth and environmental sustainability. Due to the heterogeneity of European regions, it is clear that some have significantly advanced in terms of sustainability, achieving various degrees of decoupling, whereas many others have yet to do so. This disparity must be accounted for in the analysis to determine such expenditures’ short-term and long-term effects. Our analysis also accounts for the direct impacts of policy by considering the OECD Environmental Policy Stringency (EPS) index (OECD, 2016) and mean regional political opinions as covariates.
We find heterogeneous impacts of EU structural fund investments in low-carbon economies on GHG emissions reductions across European regions, with findings highlighting the variability of these effects by regional development levels, investment periods, and emission types. In developed and transition regions, low-carbon investments significantly lowered the long-term emission trends, in both per capita and relative to GDP, across the 2007-2013 and 2014-2020 programs. However, in the least developed regions, “low-carbon” investments had a positive (and significant) impact on emissions. Additionally, the effectiveness of reducing specific gas emissions, including CO2, CH4, and N2O, exhibits variability, indicating a complex interaction between investment efficiency and pollutant type. Our analysis also discerns significant but divergent impacts on the cyclical emissions components in transition economies between the two program periods. Hence, our analysis suggests a partial inefficiency of low-carbon public investments in Europe. It calls for a refined spatial and temporal allocation of these investments.
The heterogeneous impacts of the climate action efforts derived from the ESI funds can be attributed to several critical factors. These factors include the inherent complexities of reducing emissions, the intrinsic differences among European regions in terms of capabilities and resource endowments, and the relevance of development paths, as suggested by the Kuznets environmental curve (Bataille et al., 2021; Saraji and Streimikiene, 2023). For instance, it is possible that any additional funding may increase energy consumption and GHG emissions for investments below a certain level. This is rooted in how investments are targeted and their immediate impacts, which might not always be aligned with long-term sustainability goals. In many developing regions, the primary focus of initial investments may be on establishing basic infrastructure and boosting economic output quickly to meet immediate developmental needs. Additionally, low-carbon economy investments typically require substantial up-front costs and the deployment of heavy infrastructure, which can drive significant energy consumption and emissions. The transition to sustainable energy systems includes upgrading grids, developing renewable energy sources, and improving energy resilience. Though essential in the long term, these investments can increase short-term emissions due to the intensive energy and resources required during their implementation phase. On another front, there is a risk that investments have been used inefficiently or inappropriately, for example, as a countercyclical buffer to stabilise the economy during recessions or by investing in large infrastructure projects (aka “white elephants”) that do not generate the expected impacts. Arguments like these could align with the heterogeneous effects of public investment on climate outcomes.
Therefore, the differential impacts on GHG emissions have their origin in the cyclical nature of economic activities. Recent literature offers valuable insights into the interplay between environmental impacts and the business cycle. Doda (2014) finds that CO2 emissions increase during economic upturns, exhibit greater volatility compared to GDP, and show a procyclicality that is more pronounced in higher GDP per capita economies, suggesting wealthier economies have more stable emission patterns through economic cycles. Sheldon (2017) finds an asymmetrical response of emissions to economic changes attributed to the timing of capital investments and technological advancements, indicating the significant role of corporate decisions in environmental impacts. Moreover, Klarl (2020) finds CO2 emissions’ elasticity to GDP changes to be significantly higher during recessions in the United States, showcasing the dynamic elasticity of emissions in response to economic conditions.
In addition, climate and energy policy tools have proven some degree of effectiveness, but their success often hinges on countries’ stringency and the complementarity among strategies. Diakoulaki and Mandaraka (2007) find that despite considerable efforts, the expected acceleration in decoupling post-Kyoto Protocol was not uniformly achieved. For Europe, Naqvi and Zwickl (2017) and Ivanova et al. (2017) also observe significant disparities in decoupling across countries, sectors, pollutants, and regions. Papież et al. (2022, 2021) further report mixed results in the EU’s long-term decoupling efforts, pointing to disparities between older and newer member states and emphasising the role of energy intensity and structure in emissions reductions. For OECD countries, Chen et al. (2018), focusing on the key factors affecting CO2 emissions, conclude that while reductions in energy intensity are associated with decreased CO2 emissions, economic growth inevitably tends to increase CO2 emissions. This implies that emissions reductions are hard to accomplish unless economic growth is accompanied by significant improvements in energy efficiency and a shift towards less carbon-intensive energy sources.
Based on the literature discussed above, our paper contributes to the analysis of sustainable development and economic policy in European regions. By analysing how low-carbon investments from structural funds impact emissions reductions within business cycle fluctuations, we contribute to the understanding of decoupling economic growth from environmental degradation. Our results emphasise the importance of tailored investment strategies in achieving sustainable development goals, with a special focus on less developed regions to mitigate potential increases in emissions as they grow economically.
The remainder of this study is structured as follows. Section 2 discusses the relevant strands of literature and the conceptual framework of our study and situates our contribution. Section 3 presents the data and methods. Section 4 presents the empirical results of the econometric estimations. Finally, Section 5 concludes by discussing the policy implications of our results.
2 Decoupling emissions: related literature
Our study builds on the extensive literature analysing the relationship between economic growth and GHG emissions. The Environmental Kuznets Hypothesis indicates that there is a connection between a country’s income level and the quality of its environment: it can be said that an increase in wealth generates greater demand from the population for a cleaner and healthier environment. This demand, in turn, drives governments to implement and enforce stricter environmental laws and regulations to protect the environment. While wealthier countries tend to have the resources and political will to adopt and enforce these effective environmental policies, regions with more constrained economic resources may find it more challenging to establish and maintain high environmental standards, resulting in more significant environmental challenges (O’Connor and Turnham, 1992; Doyle and Stiglitz, 2014).
Certainly, economic development can lead to improved environmental outcomes by decoupling emissions from economic output. This is central to sustainable development and indicates that an economy can grow without a proportional increase in environmental degradation. However, the process is influenced by a multitude of factors, not only including income levels but also other aspects such as consumption patterns and regional characteristics. Shuai et al. (2017) analysed the country-by-country turning points of the carbon Kuznets curve, finding a positive correlation between a country’s income level and its likelihood to fit the Kuznets hypothesis, with higher income economies achieving more quick tipping points than their poorer counterparts. While economic development level is linked to faster environmental improvement transitions, other variables also play a relevant role. Ivanova et al. (2017) analyses household consumption across 177 regions in 27 EU countries, finding wide variations in per capita GHG emissions across regions (0.6 to 6.5 tCO2e/cap in 2007). Regarding the drivers behind carbon footprints, the authors identified the relevance of household size, whether an area is urban or rural, the population’s education level, expenditure patterns, local temperature, resource availability, and the carbon intensity of the local electricity mix.
Yet a significant strand of literature doubts that the downward-sloping relationships between higher-income countries and pollution levels – which might imply a decoupling of the economic activity with emissions – is enough to face climate change efficiently. Certainly, there are concerns about the possible shift in the production of pollution-intensive goods from developed to developing countries with less stringent environmental regulations (Hertwich and Peters, 2009; Levinson, 2023). This would weaken the Kuznets hypothesis since developed regions may reduce their at-home pollution levels by offshoring the production of environmentally harmful goods to nations where environmental regulations are less strict (Fischer and Heutel, 2013). While this points out that high-income countries should also look into reducing aggregate demand as a strategy for significant emissions reductions (Knight and Schor, 2014), it also requires the implementation of agreements at different scales: regional and global (Oberthür and Ott, 1999).
A growing body of evidence indicates that such policies and international agreements are increasingly impacting emission trends, as reflected in the economic practices within countries, contributing to our understanding of the decoupling mechanism (Fischer and Heutel, 2013; Hildén et al., 2014; Haberl et al., 2020). Decoupling analyses aim to categorise the patterns of how emissions respond to economic growth – whether they increase at a slower rate, decrease, or stabilise as the economy grows (relative and absolute decoupling) or increase faster than the economy (coupling).
Cohen et al. (2018) provide valuable quantitative analysis that helps recognise how major emitting countries are progressing in decoupling emissions. Using information from OECD countries for 2001-2015, Chen et al. (2018) analyse how economic growth can be decoupled from CO2 emissions, identifying three key forms: recessive, where emissions drop due to efficiency gains or energy mix shifts, not economic growth; weak negative, where efficiency improvements do not fully offset emissions from economic expansion; and strong, where economic and population growth do not lead to higher emissions, thanks to effective policies, technological progress, and structural economic changes.
Building on this understanding of the decoupling patterns, it is imperative to analyse emissions in the context of business cycles due to their substantial implications for environmental policy and economic management, particularly regarding the relationship between economic activity, regulation, emissions and energy consumption (Fischer and Heutel, 2013). In fact, emissions tend to be procyclical, increasing during economic booms and decreasing during recessions (Heutel, 2012). A fundamental question relates to the causal relationship between economic output and energy consumption, especially targeting business cycle fluctuations (Thoma, 2004; Narayan et al., 2011). Energy consumption notably responds to changes in industrial production and macroeconomic conditions, reinforcing the concept of the procyclicality of emissions relative to GDP observed in broader research. For instance, Thoma (2004) demonstrates that changes in macroeconomic conditions cause significant changes in electrical usage, particularly in the commercial and industrial sectors.
Therefore, the link between economic growth, energy consumption, and GHG emissions is clearly established in the short term. For example, Peters et al. (2012) claim that economic downturns typically lead to a reduction in energy-intensive activities and lower CO2 emissions. Notwithstanding, the authors discuss that the 2008-2009 financial crisis saw a quick dip in GDP followed by a swift rebound in emissions by 2010, driven by falling energy prices, government economic incentives, and prior economic growth in developing countries, quickly returning to high emission levels post-crisis. Similarly, Doda (2014) finds that CO2 emissions are procyclical and more pronounced in countries with higher GDP per capita. Hence, the procyclicality also implies that economic growth without specific interventions often leads to increased emissions. Another fact is that emissions exhibit greater volatility compared to GDP, responding more dramatically to economic changes, and the extent to which emissions volatility exceeds GDP volatility decreases with the income level of the economies (Doda, 2014). This implies that more stable emission patterns through economic cycles in developed regions could be linked to more efficient energy use, a greater share of services in the economy, and potentially more effective environmental regulations and policies. In addition, emissions decrease more significantly during economic contractions than during expansions, suggesting an elastic response to economic downturns and inelasticity to economic growth (Sheldon, 2017).
Likewise, Klarl (2020) examines the fluctuation of CO2 emissions’ elasticity to GDP in the United States from 1973 to 2015, revealing that this elasticity is not constant and varies with economic conditions. He showed that emissions respond more sensitively to GDP changes during recessions than periods of economic normality, with elasticity often exceeding one in downturns. This indicates that CO2 emissions decrease more rapidly than economic output in recessions, highlighting the impact of economic cycles on environmental outcomes. The finding that emissions exhibit greater sensitivity during economic downturns suggests that the capacity of an economy to influence environmental outcomes can be significantly conditioned by its current economic state.
In a similar vein, Naqvi (2021) finds that, on average, EU regions have managed to decouple emissions from economic output, with the most significant emission reductions occurring before the 2008 financial crisis. After 2008, the trend towards decoupling weakened, with instances of re-coupling between emissions and economic output observed. The author uses the EPS index, finding that stronger environmental policies effectively reduce emissions. However, the effectiveness varies significantly across different types of emissions and is also influenced by the income levels of the regions. Similarly, with the aim of monitoring the effectiveness of the EU’s environmental policies, Delgado Rodríguez et al. (2018) developed the EU-cyclical environmental performance index by applying dynamic factor analysis over the period 1950 to 2012. This index demonstrates that the implementation of stricter emission targets, initiated particularly since the adoption of the 5th Environmental Action Programme (EAP), has effectively contributed to restraining the growth of CO2 emissions across the EU. Furthermore, the index reveals heterogeneity among the member states, distinguishing between countries that are leading in emission linkage – thus enhancing their alignment with EU environmental standards – and those lagging behind, which may require additional incentives or policy measures to improve compliance with EU regulations. In addition, Papież et al. (2022) evaluate the EU’s energy policy effectiveness in achieving long-term decoupling and reports mixed findings. The study notes progress in detaching economic growth from emissions under the EU ETS and production-based accounting in both EU-15 and new member countries. However, they find no advancement in separating economic growth from consumption-based accounting emissions in either group.
In summary, environmental policies must confront the broader economy, influenced by external macroeconomic shocks and internal technological changes. Under Kuznets’ hypothesis, decoupling might appear problematic since countries would only begin to improve their environmental performance after reaching a certain level of wealth, implying a dependency between economic growth and environmental improvement. Therefore, seeking economic development without the corresponding environmental degradation is a great challenge for any policy framework.
3 Data and methods
We focus on emission data reported by the Emissions Database for Global Atmospheric Research (EDGAR), managed by the Joint Research Centre of the European Commission. EDGAR provides quantitative estimates of global GHG emissions by country and sector, including CO2, CH4, N2O, and F-gases, using a methodology aligned with Intergovernmental Panel on Climate Change (IPCC) guidelines (Crippa et al., 2023). Specifically, we use European regional data at the NUTS 2 level of aggregation, focusing on aggregated GHG emissions and their breakdown into CO2, CH4, and N2O emissions.
Our primary objective is to analyse the impacts on GHG emissions resulting from expenditure on the thematic objective related to the low-carbon economy in the 2007-2013 and 2014-2020 ESI funds programmes (a more detailed definition of the expenditure variables is presented below). In the econometric modelling setup, we use other relevant covariates to control for economic activity, sectoral composition, and local policy stringency that influence emissions across European regions.
3.1 Data
ESI funds and eligibility
We use data from the Cohesion Open Data Platform, which provides a granular view of the thematic content financed during the 2007-2013 and 2014-2020 programmes. The datasets for both periods encompass cumulative categorisation data on projects as reported in the European Regional Development Fund (ERDF), Cohesion Fund (CF), and European Social Fund (ESF) programme final implementation reports as presented in the closure documentation.222For 2007-2013 raw categorisation data see: https://cohesiondata.ec.europa.eu/2007-2013-Categorisation/2007-2013-Cohesion-categorisation-raw-data-from-cl/qt3c-wiyg/about_data.
For 2014-2020 raw categorisation data see: https://cohesiondata.ec.europa.eu/2014-2020-Categorisation/ESIF-2014-2020-ERDF-CF-ESF-raw-categorisation/xe4p-7b9q/about_data. These data contain the volume of EU support allocated to projects alongside a combination of categorisation dimension codes, including priority themes, the form of finance, the territorial context, location (using the Nomenclature of Territorial Units for Statistics, NUTS), and the economic dimension. We aggregate investments at NUTS level 2 (NUTS 2), as defined in version 2021. To do this, we consider all related investments, regardless of their original level of geographic specificity in the original database – be it national, NUTS-1, or even more granular at NUTS-3.
Figure 1 shows the calculated distribution of total ESI funds expenditures reported by region for the periods 2007-2013 and 2014-2020, along with the eligibility of the regions to access these funds. Eligibility is based on an income classification of regions, categorising them into three groups: LDR (less developed regions), TER (transition regions), and DER (developed regions). In all econometric exercises in this paper, we use this regional income categorisation fixed.333The details of the eligibility classification, according to the ESI funds assigned in 2014-2020, can be found here: https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=CELEX:32014D0099.
The total expenditures reflect the targeted efforts to reduce economic disparities within the EU. Certain regions consistently receive higher funds according to ongoing economic development needs or projects requiring sustained investment. The disparities in the distribution of funds consistently with eligibility show that less developed regions in Southern and Eastern Europe generally receive higher per capita funding than Western and Northern Europe regions.
There are also variations in the distribution of funds between programming periods, which must be explained by changes in policy focus. Clearly, the EU’s structural funds programmes significantly evolved from their initial focus on procuring the catching-up of less developed regions towards a more strategic approach. In the 2007-2013 programming period, a transition towards integrating environmental and climatic considerations began, emphasising energy efficiency, renewable energies, and sustainable resource management. However, the transition to a more strategic approach was cemented during the 2014-2020 programming period, with a clear directive to channel funds towards smart, sustainable, and inclusive growth.
The 2014-2020 funding period introduced thematic objectives (TOs), including: Research and Innovation (TO-1), Information and Communication Technology (TO-2), Competitiveness of Small and Medium-Sized Enterprises (TO-3), Low-Carbon Economy (TO-4), Climate Change Adaptation, Risk Prevention, and Management (TO-5), Environmental Protection and Resource Efficiency (TO-6), Sustainable Transport and Removing Bottlenecks in Key Network Infrastructures (TO-7), Employment and Labour Mobility (TO-8), Social Inclusion and Combating Poverty (TO-9), Education, Training, and Vocational Training for Skills and Lifelong Learning (TO-10), and Institutional Capacity Building and Efficient Public Administration (TO-11).
Low-carbon economy funds
We define as “climate action funds” all expenditures aimed at reducing emissions in both programming periods. Despite thematic changes between the two periods, the raw categorisation data allows for the creation of a comparative set of expenditure objectives to ensure comparability and coherence in our analysis.
For the 2014-2020 period, the low-carbon expenditures are straightforward to determine since there is a categorical variable related to thematic objective 4 (TO-4), labelled as “low-carbon economy”. The aim of the expenditure can be tracked by the reported “intervention field” in the data. Specifically, related to TO-4, the intervention fields include renewable energy (wind/solar/biomass), energy efficiency renovation of public infrastructure, cycle tracks and footpaths, among others (see Table A1 in the Appendix).
For the 2007-2013 program, there is no categorical variable explicitly associated with the low-carbon economy as in the 2014-2020 program. However, the raw categorisation data describes the “priority” related to the reported expenditures, serving as a proxy for the aim of the expenditure. We use this information to create a low-carbon expenditures variable for this period. Priority labels related to this period include cycle tracks, intelligent transport systems, renewable energy (wind/solar/biomass/hydroelectric), energy efficiency, co-generation, energy management, and promotion of clean urban transport. The priority codes are reported in Table A1 of the Appendix.
A critical aspect of the ESI funds expenditure data is the absence of detailed annual expenditure information at the regional level. While reported expenditures can be easily tracked at the regional level, the reports correspond to the entire programming period. To our knowledge, there is no disaggregated, time-varying consolidated official information. Nevertheless, the cohesion data platform presents estimations of the expenditure by NUTS 2, which are aggregated for the most significant types of structural funds only. Therefore, to address our need for yearly data, we have opted for two approaches: one assumes that expenditures are distributed evenly over time (dividing the reported expenditures over the programming period), and the other assumes they follow the same pattern as the total aggregated expenditure reported by the cohesion data platform.444Data available at: https://cohesiondata.ec.europa.eu/Other/Historic-EU-payments-regionalised-and-modelled/tc55-7ysv/about_data. Since both strategies yielded qualitatively similar conclusions, we present in the main text the results obtained with the first approach, while those derived from the second are included in the Appendix as a robustness check (Tables A2 and A3).
Supplementary covariates
Economic variables were obtained from the ARDECO (the Annual Regional Database of the European Commission’s Directorate General for Regional and Urban Policy), which contains annual data at the regional level and is maintained and updated by the Joint Research Centre of the European Commission. We use variables such as the real gross domestic product (GDP), the population, and the real gross value added (GVA) for all industry sectors according to NACE Rev. 2 definitions. These sectors are A: Agriculture, Forestry and Fishing; B–E: Industry; F: Construction; G–J: Wholesale, Retail, Transport, Accommodation & Food Services, Information and Communication; K–N: Financial & Business Services; and O–U: Non-market Services. The choice of these variables is motivated by their relevance to understanding how economic activity is related to emissions. In particular, we include industry-level GVA to control direct sources of emissions and to better capture regional emission profiles, given that the participation of industries varies significantly across European regions.
To assess the effects of environmental policies on emissions, we use the EPS index (OECD, 2016). This index, developed by Brunel and Levinson (2013) and further refined by Botta and Koźluk (2014), provides a standardised composite measure incorporating both market and non-market-based policies, covering 27 OECD countries over the period from 1990 to 2020.
Finally, to determine the political ideology at the regional level, we calculate the regional weighted average of political ideologies using the results of European subnational election data. To measure political ideology, we use the “left/right” dimension variable reported by ParlGov (Döring et al., 2022) and complemented by CHES-Europe (Jolly et al., 2022), which takes values in the interval 0–10, with values close to zero representing extreme left ideology and values close to 10 representing extreme right ideology. We use EU-NED (Schraff et al., 2023), which reports sub-national election data across European countries over the last three decades at the party level. We assign each party at the regional level an ideology by combining information from ParlGov and CHES-Europe. These databases are integrated within the Party Facts platform, which facilitates the incorporation of party-level data.555The data is available at: https://www.chesdata.eu/ches-europe and https://www.parlgov.org/. Also see: https://partyfacts.herokuapp.com/. We include this variable to provide additional controls for political ideology. While our focus is not on analysing how politics affects emissions, this is a growing area of interest in political science. For instance, Jahn (2021) highlights regional variations in the impact of populist governance on environmental policies, showing that right-wing populist administrations in North Western and Eastern Europe are linked to heightened GHG emissions, whereas their left-wing counterparts in Southern Europe are associated with decreases in emissions.
Distribution of emissions and low-carbon economy funds
Figure 2 presents a set of maps illustrating different facets of the environmental impact and distribution of the Cohesion Funds. In the upper part, three maps show average GHG emissions in shades of blue to red, where blue denotes lower levels and red denotes higher levels. The first map, on the left, represents the average emissions for 2015-2019 in teragrams of CO2 equivalent per year (Tg CO2eq/year). The other two maps compare the average annual per capita emissions for 2000-2004 and 2015-2019, respectively. These maps show that, on average, towards the end of the 2010s, most European regions emitted less than 50 teragrams of CO2. However, there is significant heterogeneity in per capita emissions, highlighting areas of intense economic activity or low population density. Comparing the early 2000s and late 2010s, it is evident that many regions with initially high emissions tended to reduce them (see, for instance, the Castile region in Spain), while some regions maintained or increased their per capita emission levels (for instance, the Masovian region in Poland).
The lower maps in Figure 2 show the calculated funds allocated to promote a low-carbon economy during 2007-2013 and 2014-2020 programmes, with the amounts received per capita in colour scales ranging from less than 100 euros to almost 1,000 euros per person. In both periods, most of the funds went to LDRs, with significant differences in the allocated amount, in agreement with the results presented in Figure 1.
This pattern raises questions about the effectiveness of the allocation of funds relative to emission levels. In Figure 3, we plot the total GHG emissions per capita versus the total expenditure in low-carbon economy funds per capita. Emissions are considered at the beginning of the corresponding programming period. The plot exhibits great moderation in emission levels per capita for most European regions and shows that higher funds per capita were consistently allocated to less developed regions. However, a positive correlation pattern in the allocation is less clearly seen in both panels, suggesting that these funds are not necessarily allocated to the regions with the highest need to reduce emissions.
As a preliminary analysis, we estimate the nexus between expenditure on low-carbon economy funds and total GHG emissions using the following regional-level model:
where is an emission indicator for region at time , and and are the cumulative yearly programme expenditures per capita (in log) for the periods 2007-2013 and 2014-2020, respectively, with denoting lags to acknowledge potential delayed effects. We include fixed effects for location () and time () to control for time-invariant unobserved heterogeneity across regions as well as uniform temporal shocks, such as those associated with environmental directives instituted at the European Union level. Finally, the term is the estimation residual.
In the left-hand in Eq. (1), the term serves as a reference notation for different emission variables. Specifically, we use four different dependent variables based on the total emissions of region at time . These are the levels of emissions per capita (in log), defined as GHGEpci,t=; the emission intensity relative to the GDP, defined as GHGEii,t=; and the growth rates of GHGEpci,t and GHGEii,t.
The estimation results are presented in Table 1. We find that expenditures have a statistically significant positive relationship with per capita emissions, as seen in models 1 and 2, indicating a perverse effect against the intended impact of such investments. When examining the growth rates of emissions per capita, a negative relationship appears with the three-year lag in spending (model 3). However, in model 4, which accounts for a five-year lag, this relationship is not statistically significant. Regarding emission intensity, we find that low-carbon expenditures are linked to decreases in this measure, as demonstrated in models 5 and 6. At the same time, the growth rate of emission intensity does not show a significant effect related to the expenditures.
| GHGEpc | GHGEi | |||||||
| Levels (log) | Growth rates | Levels (log) | Growth rates | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | |
| Low Carbon Exp. pc (log) | 0.004 | 0.081∗∗ | 0.130∗∗ | 0.042 | ||||
| (0.063) | (0.036) | (0.064) | (0.038) | |||||
| Low Carbon Exp. pc (log) | 1.169∗∗∗ | 0.008 | 0.438∗∗∗ | 0.009 | ||||
| (0.091) | (0.053) | (0.093) | (0.055) | |||||
| Low Carbon Exp. pc (log) | 0.149∗∗ | 0.052 | 0.123∗∗ | 0.028 | ||||
| (0.062) | (0.035) | (0.063) | (0.037) | |||||
| Low Carbon Exp. pc (log) | 1.633∗∗∗ | 0.067 | 0.601∗∗∗ | 0.091 | ||||
| (0.140) | (0.080) | (0.142) | (0.084) | |||||
| Observations | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 |
| Adjusted R2 | 0.382 | 0.375 | 0.179 | 0.178 | 0.580 | 0.580 | 0.039 | 0.039 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||||
This preliminary exercise demonstrates that the relationship between low-carbon expenditures and GHG levels, as well as the growth rates of emissions per capita (GHGEpc) and emission intensity (GHGEi), can be mixed. These mixed results suggest that the econometric specification might be masking underlying factors that could restrict the effectiveness of low-carbon funds. One relevant limiting factor is the persistent disparities among European regions. Indeed, the development level of regions is a key factor influencing the effectiveness of cohesion funds (Boldrin and Canova, 2001; Rodríguez-Pose and Fratesi, 2004). However, many aspects related to the heterogeneity of European regions can explain the mixed results. For instance, less developed regions might use these funds to build sustainable infrastructure to reduce emissions, while more developed regions, having already made progress in emissions efficiency, might experience a more moderate impact from these funds. Therefore, the empirical strategy must consider that short-term outcomes may differ from longer-term impacts. This consideration is crucial for environmental policy because it suggests that sustainability investments may be misallocated, leading to unintended or unfavourable effects on emissions.
3.2 Methodology
The research on decoupling emissions from economic activity primarily hinges on analysing the emissions-GDP elasticity, which measures how emissions change with GDP and helps understand the link between economic growth and environmental impact. The central model is as follows:
where represents the log of total emissions at time , denotes the log of real GDP, denotes these first differences operator, and represents the error term. The parameters and capture the intercept and the elasticity of emissions with respect to GDP, respectively.
However, as Cohen et al. (2018) pointed out, this approach has some important caveats. Looking at the growth of emissions in relation to real GDP growth for the world’s 20 largest emitters, the authors showed a positive emissions-output elasticity across all countries, suggesting a strong correlation between economic activity and emissions. However, by decomposing emissions and real GDP into their trend and cyclical components, they show clearer evidence of decoupling in richer nations, particularly in European countries. The trend components of this analysis, which focus on long-term movements rather than short-term fluctuations, indicate that while cyclical components are still closely tied to GDP growth, the trend components are not. For several countries, including Italy, the United Kingdom, and Germany, the trend elasticities for emissions are either close to zero or negative, suggesting that the long-term component of emissions growth has decoupled from GDP growth (Cohen et al., 2018).
Our methodology shifts from this type of analysis, which primarily assesses the impact of GDP growth on emissions, to a comparative panel data analysis focusing on the effects of expenditure in the low-carbon economy. We specifically highlight the role of cohesion fund investments in fostering low-carbon economies to curb emissions. Therefore, it is central in our methodology to decompose time series variables, such as emissions, GDP, and GVA, into their cyclical and trend components for each region independently. Then we consider these time series in two different panel data model frameworks: regional trends and regional cycles of emissions. In this way, we analyse the difference between long-term trends and cyclical effects of the impact of low-carbon economy funds on GHG emissions.
In the first phase of our methodology, we apply the Hodrick-Prescott (HP) filter to each region’s emissions per capita and emissions intensity time series to separate them into their corresponding cyclical and trend components. This method allows us to distinguish short-term fluctuations associated with the economic cycle from the long-term trajectory that reflects structural changes and persistent trends in emissions. Therefore, the total emissions time series of the region , , is decomposed in its trend component and its cycle component .
After decomposing the GHGEpc and GHGEi time series, we independently analyse the panels of regional emission trends and cycles, considering the regional development level differences. The model for emissions trends is defined as follows:
where the superscript specifies that we conduct the estimation on time series trend components. This way, is the trend of the logarithm of emissions per capita for NUTS 2 region at time , similarly is the trend component of the GDP (in log) of region at time , and and are the yearly cumulative programme expenditures (in log) at time . We include the development level of region using the dummy variable , where . The term consider additional control variables such as the dummy variables capturing the effect of stronger stringency of environmental policies, denoted by 666Following Naqvi (2021), we measure the impact of stronger environmental policies employing a dummy variable that takes the value of one when EPS index is greater or equal than 2.5., and the term , which indicates the average political ideology of the region, determined by the outcomes of local elections (see data section for details). This model incorporates fixed effects for both location () and time (), removing biases related to unobserved fixed regional heterogeneity and factors that vary over time but impact all regions equally.
Additionally to Eq. (3) focusing on the trend of regional emissions, we also consider the impact of low-carbon investments on the cyclical components of emissions using the following specification:
Note that the superscript now represents the emissions’ cyclical components, therefore, is the cycle component of the GDP (in log) of region at time .
In our specifications, Eq. (3) and (4), the parameters of interest are , , and , which capture how a region’s development level can influence structural funds’ effect in emissions components. The aim is to explore whether the impact of these expenditures varies according to the economic development level of the region, affecting the cyclical and trend components of emissions.
As a robust check, in our specifications, we replace GDP with the set of GVA variables for all economic sectors. This allows us to better control how regional productive structures affect GHG emissions.
4 Results
4.1 Effect on emissions trend
Table 2 presents the estimation results of Eq (3) using as the dependent variable the trends for GHGEpc or GHGEi. The results related to the expenditures for the different groups of regions are presented with respect to the most developed regions (DER). The table reveals mixed patterns that underline the importance of regional development level in the influence of expenditure.
| Dependent variable: | ||||||
| GHGEpc | GHGEi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.158∗∗ | 0.138∗∗ | 0.012 | 0.159∗∗ | 0.171∗∗∗ | 0.108∗∗ |
| (0.069) | (0.057) | (0.054) | (0.062) | (0.053) | (0.050) | |
| Low Carbon Exp. pc | 0.558∗∗ | 0.817∗∗ | 0.049 | 0.806∗∗∗ | 1.321∗∗∗ | 0.802∗∗ |
| (0.226) | (0.415) | (0.398) | (0.204) | (0.389) | (0.371) | |
| Low Carbon Exp. pcTER | 0.029 | 0.111 | 0.127 | 0.111 | 0.059 | 0.004 |
| (0.132) | (0.116) | (0.110) | (0.119) | (0.108) | (0.103) | |
| Low Carbon Exp. pcTER | 2.156∗∗∗ | 1.177 | 3.539∗∗∗ | 2.539∗∗∗ | 1.974∗∗ | 2.895∗∗∗ |
| (0.444) | (0.818) | (0.784) | (0.400) | (0.766) | (0.732) | |
| Low Carbon Exp. pcLDR | 0.730∗∗∗ | 0.620∗∗∗ | 0.489∗∗∗ | 0.361∗∗∗ | 0.356∗∗∗ | 0.315∗∗∗ |
| (0.085) | (0.071) | (0.069) | (0.077) | (0.067) | (0.064) | |
| Low Carbon Exp. pcLDR | 2.470∗∗∗ | 3.329∗∗∗ | 2.468∗∗∗ | 1.382∗∗∗ | 2.154∗∗∗ | 1.666∗∗∗ |
| (0.226) | (0.427) | (0.408) | (0.204) | (0.399) | (0.381) | |
| GDP | 0.209∗∗∗ | 0.297∗∗∗ | 0.631∗∗∗ | 0.525∗∗∗ | ||
| (0.017) | (0.019) | (0.015) | (0.018) | |||
| EPS | 0.009∗∗ | 0.011∗∗ | 0.005 | 0.010∗∗ | ||
| (0.005) | (0.004) | (0.004) | (0.004) | |||
| Political Ideology | 0.016∗∗∗ | 0.001 | 0.002 | 0.007∗∗ | ||
| (0.004) | (0.004) | (0.003) | (0.003) | |||
| GVAA | 0.044∗∗∗ | 0.074∗∗∗ | ||||
| (0.013) | (0.012) | |||||
| GVAB-E | 0.210∗∗∗ | 0.089∗∗∗ | ||||
| (0.015) | (0.014) | |||||
| GVAF | 0.016 | 0.095∗∗∗ | ||||
| (0.014) | (0.013) | |||||
| GVAG-J | 0.0001 | 0.111∗∗∗ | ||||
| (0.029) | (0.027) | |||||
| GVAK-N | 0.327∗∗∗ | 0.506∗∗∗ | ||||
| (0.027) | (0.025) | |||||
| GVAO-U | 0.324∗∗∗ | 0.330∗∗∗ | ||||
| (0.032) | (0.030) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.534 | 0.626 | 0.664 | 0.791 | 0.798 | 0.820 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. | ||||||
Therefore, for developed regions, the reference group, the effect of expenditures in both periods indicates a negative and significant impact in all specifications, except in model 3, where the effect is not significant for both periods. Similarly, the effect of the interaction term for transition regions is negative in all specifications, except in model 2, being non-significant for expenditures associated with the period 2007-2013 and significant for expenditures in the period 2014-2020. This implies that the net impact for transition regions is negative, meaning that, for both transition and developed regions, the long-term effect of expenditures is expected to result in a decrease in emissions, both per capita and in relation to GDP.
In contrast, a significant positive effect is observed for less developed regions, reversing the trend seen in the other development groups. This is because the estimates of the expenditures interaction terms in developing regions are not only positive and significant in all specifications but also cancel any negative effect estimated for the reference group. While explaining the mechanism behind this perverse outcome is far from the scope of this paper, this could indicate that, in less developed regions, expenditures have effectively driven economic development, albeit possibly at the cost of an increase in emissions in the long term. It is important to note that as these regions are developing, the increase in emissions resulting from factors like industrialisation, which is often carbon-intensive, cannot be ignored.
Regarding the control variables, GDP is positively related to GHGEpc and GHGEi, implying that economic growth can lead to an increase in per capita emissions in the long term and to a more efficient economy in terms of emissions relative to GDP.
The EPS dummy has a small but positive impact on per capita emissions and negative emission intensity. Although this positive impact is unexpected, it may also indicate that the effect of more stringent policies may be observed thereafter. Indeed, Naqvi (2021) analysed the cumulative effects of EPS dummy on different emission components, showing that while CH4 and N2O emissions responded strongly to policies and declined continuously across regions, CO2 emissions declined two years after implementing the policies. In the Appendix, we show that the EPS dummy has a negative impact on CH4 and N2O while positive for CO2, the latter being the main component of the total GHG emissions used in this paper (Tables A4 to A6).
The effect of the average political ideology is not robust; while it is found positive for GHGEpc, except in model 3, for GHGEi, it is positive and not significant in model 5 and negative and significant in model 6. These mixed results connect to the complex relationship between the effectiveness of sustainable policies and political ideologies. In this respect, for instance, Garmann (2014) finds that in 19 OECD countries from 1992 to 2008, right-wing governments were associated with fewer emission reductions than the centre and left-wing governments. However, he also observes that emissions were higher in governments with more participating parties, although the distinction between majority and minority governments did not significantly influence emissions.
The effect related to the gross value added by sectors reveals a differentiated impact on GHGEpc and GHGEi. For instance, the GVA of the agriculture sector, GVAA, has a negative impact on both GHGEpc and GHGEi, which suggests that the trend of the agricultural sector production contributes to a reduction of the emissions trends. While the GVA of the manufacturing industry, GVAB-E has a significant positive effect on GHGEpc and a significant negative effect on GHGEi, indicating that greater production in the manufacturing industry implies higher emissions per capita but lower emission intensity.
4.2 Effect in the emissions cycle
Table 3 presents the results of the regressions for the cyclical part of per capita emissions and emissions intensity. The results show mixed effects associated with the expenditures of both periods and the development level interacted variables. These effects are presented in reference to the most developed regions. Therefore, for developed regions, we find that the effect of expenditure in 2007-2013 is negative, although only significant in model 1. The effect of expenditure in 2014-2020 is not significant, but in specifications (2-3 and 5-6), the estimated effect is positive. These results hint that expenditure in the low-carbon economy has no robust effect on the emission cycle for the most developed regions.
| Dependent variable: | ||||||
| GHGEpc | GHGEi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.081∗ | 0.027 | 0.034 | 0.068 | 0.025 | 0.022 |
| (0.045) | (0.045) | (0.045) | (0.045) | (0.045) | (0.046) | |
| Low Carbon Exp. pc | 0.143 | 0.291 | 0.237 | 0.133 | 0.273 | 0.159 |
| (0.147) | (0.331) | (0.330) | (0.146) | (0.331) | (0.335) | |
| Low Carbon Exp. pcTER | 0.120 | 0.279∗∗∗ | 0.323∗∗∗ | 0.113 | 0.282∗∗∗ | 0.326∗∗∗ |
| (0.086) | (0.090) | (0.090) | (0.085) | (0.090) | (0.091) | |
| Low Carbon Exp. pcTER | 0.247 | 1.620∗∗ | 1.807∗∗∗ | 0.266 | 1.746∗∗∗ | 1.558∗∗ |
| (0.285) | (0.638) | (0.637) | (0.283) | (0.637) | (0.645) | |
| Low Carbon Exp. pcLDR | 0.010 | 0.003 | 0.012 | 0.018 | 0.018 | 0.016 |
| (0.056) | (0.057) | (0.057) | (0.055) | (0.057) | (0.058) | |
| Low Carbon Exp. pcLDR | 0.029 | 0.473 | 0.481 | 0.035 | 0.467 | 0.253 |
| (0.148) | (0.341) | (0.341) | (0.147) | (0.341) | (0.345) | |
| GDP | 0.458∗∗∗ | 0.519∗∗∗ | 0.538∗∗∗ | 0.463∗∗∗ | ||
| (0.028) | (0.035) | (0.028) | (0.035) | |||
| EPS | 0.025∗∗∗ | 0.030∗∗∗ | 0.017∗∗∗ | 0.016∗∗∗ | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| Political Ideology | 0.003 | 0.002 | 0.001 | 0.0001 | ||
| (0.003) | (0.003) | (0.003) | (0.003) | |||
| GVAA | 0.011 | 0.006 | ||||
| (0.008) | (0.008) | |||||
| GVAB-E | 0.112∗∗∗ | 0.082∗∗∗ | ||||
| (0.016) | (0.016) | |||||
| GVAF | 0.086∗∗∗ | 0.030∗∗ | ||||
| (0.013) | (0.013) | |||||
| GVAG-J | 0.058∗∗ | 0.199∗∗∗ | ||||
| (0.025) | (0.025) | |||||
| GVAK-N | 0.069∗∗ | 0.143∗∗∗ | ||||
| (0.033) | (0.033) | |||||
| GVAO-U | 0.187∗∗∗ | 0.0003 | ||||
| (0.037) | (0.037) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| R2 | 0.248 | 0.295 | 0.302 | 0.166 | 0.150 | 0.136 |
| Adjusted R2 | 0.193 | 0.235 | 0.241 | 0.105 | 0.078 | 0.060 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
For transition regions, the estimated effect is relatively clearer; nonetheless, determining the net effect related to each program has opposite signs. The interaction with expenditures in the 2007-2013 programme shows a positive and significant effect (except in models 1 and 4) even after discounting the effect of the reference group. The net effect of expenditure in this period suggests an increase in the intensity of emission cycles, both per capita and concerning GDP. In contrast, the effect associated with expenditure in the 2014-2020 programme is consistently negative and significant (except in models 1 and 4). Then, considering the effect of the reference group, the net effect remains negative and comparatively much stronger. Although the estimated effects for both periods are opposing, the expected effect of these expenditures is to reduce the intensity of emission cycles. For less developed regions, the effect of expenditures associated with emission cycles is negative in almost all specifications for both periods, although not significant. Thus, considering the estimated effects in the reference group, the effect of expenditures on emission cycles in less developed regions is not significant.
Regarding the other control variables, the GDP cycle has a strong positive impact on the per capita emission cycle and a negative impact on the emission intensity cycle, indicating that economic growth cycles are associated with higher per capita emissions but with a lower intensity of emissions intensity. The country-level EPS dummy shows a significant negative effect on emissions, both per capita and intensity, suggesting that stricter environmental policies are associated with lower emissions in their cycles. The region’s weighted average political ideology does not show a significant impact on emission cycles, indicating that local political outcomes, whether left or right, may not have a direct relationship with the cyclical variability of emissions.
Finally, the sectoral gross value added coefficients indicate that different sectors have varying impacts on emission cycles. For example, the manufacturing sector (B-E) has a significant positive impact on per capita emissions and a negative impact on emission intensity. Therefore, most industry sectors have a significant influence on the cyclical variations in emissions.
4.3 Effect by type of gases
We break down emissions by specific gases, considering CO2, CH4, and N2O, to assess low-carbon investments’ impacts, using the same methodology. Detailed numerical result tables can be found in the Appendix (Tables A4 to A9), allowing us to concentrate on qualitative insights.
Disaggregated GHG emissions data at the regional level allow for a better understanding of the heterogeneous impacts of the investments and can help identify specific expenditure gaps to improve policy responses. While CO2 contributes significantly to GHG emissions and changes in CO2 emissions primarily explain the impact of low-carbon funding, it is essential to consider a broader range of gases, which have different sources and economic activity patterns. In addition, decomposing GHG emissions into leading pollutant gases is relevant as European regions vary significantly in emission types, which are linked to their socioeconomic composition (Crippa et al., 2023). CO2 results from burning fossil fuels, chemical processes like cement production, and natural emissions from trees and biological materials. CH4 is emitted during fossil fuel extraction and transportation and significantly from agricultural activities such as livestock and organic waste decomposition in landfills. N2O is produced by agricultural and industrial activities, fossil fuel combustion, and wastewater treatment.
The fixed effects estimations results on CO2, CH4, and N2O emissions are presented on Tables A4 to A6 in the Appendix.
For CO2 emissions, low-carbon expenditures for the 2007-2013 and 2014-2020 programmes show significant negative impacts on emissions per capita and relative to GDP in developed and transition regions, indicating effective reductions. However, in less developed regions, these expenditures result in positive significant impacts, suggesting an increase in emissions or less effective reductions. These results are similar to those observed for overall GHG emissions (Table 2).
For CH4 emissions, low-carbon expenditures for the 2007-2013 programme generally show negative but non-significant impacts across most specifications (including emissions per capita and relative to GDP). For the 2014-2020 programme, there are significant negative impacts in transition regions but positive or non-significant impacts in less developed regions, indicating inconsistent effectiveness. Regarding N2O emissions, low-carbon expenditures for both the 2007-2013 and 2014-2020 programmes have negative impacts but are not significant in developed regions. In transition regions, the impacts are negative and significant, while in less developed regions, the impacts are positive and significant, indicating less effective reductions or increases in emissions.
These results show that low-carbon expenditures have varying impacts depending on the type of gas and the level of regional development. In developed and transition regions, these investments are generally effective in reducing CO2 emissions but exhibit mixed and often non-significant results for CH4 and N2O. In less developed regions, low-carbon expenditures frequently result in increased emissions or less effective reductions across all gases.
Tables A7 to A9 in the Appendix present the fixed effect estimation results on the cyclical components of CO2, CH4, and N2O emissions, respectively. For CO2 emissions, low-carbon expenditures for the 2007-2013 and 2014-2020 programmes show generally non-significant impacts on the cyclical component of emissions per capita and relative to GDP. There are mixed results in transition regions, with significant positive impacts for 2007-2013 expenditures and significant negative impacts for 2014-2020 expenditures. In less developed regions, the impacts are primarily non-significant, with some negative impacts on the cyclical emission component for 2014-2020 expenditures in certain specifications.
For CH4 emissions, low-carbon expenditures for the 2007-2013 programme show consistently significant negative impacts for all regions, indicating effective reductions on the cyclical emissions component. The 2014-2020 expenditures yield mixed results, with generally non-significant impacts except for significant negative impacts in transition regions. Regarding N2O emissions, low-carbon expenditures for both the 2007-2013 and 2014-2020 programmes typically show non-significant impacts. In transition regions, there are some negative differential impacts for 2014-2020 expenditures, but these are not consistently significant. In less developed regions, the differential impacts are mixed, with positive and significant for the 2007-2013 period and non-significant for 2014-2020.
The estimations for the cyclical emissions components reveal that low-carbon expenditures have mixed results, which vary depending on the type of gas used and the level of regional development. This is observed for both per capita and per unit of GDP.
5 Concluding remarks
This paper examines the impact of climate action through one of the key vehicles for European regions: the “Low-Carbon Economy” funds of the EU Cohesion and Structural Policy for the 2007-2013 and 2014-2020 programmes. Our analysis reveals that the impact of spending on low-carbon economies in reducing GHG emissions varies significantly depending on the level of regional development, according to the EU classification that distinguishes between less developed regions, regions in transition, and developed regions. We observe that, for both analysed programmes, investments in low-carbon economies are associated with a statistically significant decrease in the long-term trend of total emissions, both per capita and relative to GDP, in developed and transition regions. However, in less developed regions, these investments have a statistically significant positive impact, indicating either an increase in emissions or a less effective reduction. Furthermore, estimations based on specific gas types also demonstrate differential impacts.
These results indicate a clear dichotomy: investments yield significant reductions in long-term emissions trends in developed and transition regions but appear to have the opposite effect in less developed regions. This aligns with insights from Doda (2014) and Sheldon (2017), who emphasise the variability of emission responses to economic cycles, particularly noting the stabilising effect of wealthier economies on emission patterns. Conversely, in less developed regions, the counterintuitive increase in emissions suggests a subtle interplay between economic development and environmental sustainability efforts, reminiscent of the disparities identified in the existing literature (Naqvi and Zwickl, 2017; Ivanova et al., 2017).
In addition, our analysis showed that European regions’ responses to low-carbon economies funds vary significantly among programmes, emphasising the importance of taking a nuanced approach when evaluating the impact of these investments on environmental sustainability. This involves considering the particularities of each program, the intrinsic differences between regions and types of emissions.
Our findings support the idea of a refined approach to the implementation of climate policy objectives through the EU cohesion policy. We agree with other studies claiming that a diverse set of guidelines must be implemented to effectively reduce emissions rather than relying on individual initiatives (Brinkley, 2014). Consequently, economy-wide assessments, sector-specific trends, planning expertise in implementing place-based approaches, and extensive case studies can help underscore the complexity of decoupling economic growth from pollutant emissions (Brinkley, 2014; Naqvi and Zwickl, 2017).
In general, the need for differentiated policies to achieve economic development is an issue that has been widely studied since the creation of the Cohesion Policy for European regions, underscoring the importance of addressing persistent inequalities between EU regions. Very early on, Boldrin and Canova (2001) highlighted that these policies tend to be more redistributive than growth-oriented, reflecting political compromises over economic efficiency. However, investments in education and human capital, which constitute a significant portion of these funds, have shown medium-term positive impacts, emphasising the critical role of human capital development in fostering long-term regional growth and convergence (Rodríguez-Pose and Fratesi, 2004). Moreover, Rodríguez-Pose and Garcilazo (2015) and Charron et al. (2014) argue that the quality of local and regional governance directly affects economic growth and the effectiveness of structural funds. They find significant variations in governance quality within countries, such as levels of corruption and efficiency of public services, which have substantial implications for policy design. Iammarino et al. (2019) emphasise the need for tailored policies that are sensitive to the structural prospects of different regions, aligning with the Barca report’s advocacy for place-based strategies (Barca, 2009). These insights align with the broader literature on the Environmental Kuznets Hypothesis, which suggests that economic development can lead to improved environmental outcomes, but only if accompanied by effective governance and tailored policy interventions.
Our research underscores the pressing need for tailored environmental policies to optimise interventions across European regions. The differential impacts of cohesion policy investments on regional emission trends and cycles highlight the necessity of moving away from one-size-fits-all solutions towards development-sensitive and emission-specific strategies. Special attention should be given to less developed regions, which, although not major emitters per capita currently, risk becoming significant pollution producers as they develop economically. Aligning investment strategies with regional profiles can enhance the effectiveness of emissions reduction efforts, contributing to a more nuanced policy dialogue on sustainable development. This underscores the critical intersection of regional growth, sustainability investments, and environmental outcomes within the EU’s policy framework.
Acknowledgements
The authors acknowledge the financial support of the DECIPHER (Decision-making framework and processes for holistic evaluation of environmental and climate policies) project, funded by the Horizon Europe: Research and Innovation programme under grant agreement No. 101056898.
Appendix
Low-carbon economy definitions by ESI programmes
| 2014-2020 funding program | |
| Intervention field code | Intervention field description |
| 009 | Renewable energy: wind |
| 010 | Renewable energy: solar |
| 011 | Renewable energy: biomass |
| 012 | Other renewable energy (including hydroelectric, geothermal and marine energy) and renewable energy integration (including storage, power to gas and renewable hydrogen infrastructure) |
| 013 | Energy efficiency renovation of public infrastructure, demonstration projects and supporting measures |
| 014 | Energy efficiency renovation of existing housing stock, demonstration projects and supporting measures |
| 015 | Intelligent Energy Distribution Systems at medium and low voltage levels (including smart grids and ICT systems) |
| 016 | High efficiency co-generation and district heating |
| 023 | Environmental measures aimed at reducing and / or avoiding greenhouse gas emissions (including treatment and storage of methane gas and composting) |
| 043 | Clean urban transport infrastructure and promotion (including equipment and rolling stock) |
| 068 | Energy efficiency and demonstration projects in SMEs and supporting measures |
| 070 | Promotion of energy efficiency in large enterprises |
| 090 | Cycle tracks and footpaths |
| 2007-2013 funding program | |
| Priority code | Priority description |
| 06 | Assistance to small and medium enterprises (SMEs) for the promotion of environmentally-friendly products and production processes (…) |
| 24 | Cycle tracks |
| 28 | Intelligent transport systems |
| 39 | Renewable energy: wind |
| 40 | Renewable energy: solar |
| 41 | Renewable energy: biomass |
| 42 | Renewable energy: hydroelectric, geothermal and other |
| 43 | Energy efficiency, co-generation, energy management |
| 25-52 | Promotion of clean urban transport |
Estimation results modelled expenditures
| Dependent variable: | ||||||
| GHGEpc | GHGEi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.182∗ | 0.167∗ | 0.004 | 0.185∗∗ | 0.213∗∗∗ | 0.134∗ |
| (0.105) | (0.085) | (0.081) | (0.093) | (0.079) | (0.075) | |
| Low Carbon Exp. pc | 1.239∗ | 5.703 | 2.736 | 1.786∗∗∗ | 9.205∗∗ | 3.117 |
| (0.715) | (3.972) | (3.798) | (0.635) | (3.684) | (3.512) | |
| Low Carbon Exp. pcTER | 0.051 | 0.151 | 0.178 | 0.198 | 0.107 | 0.046 |
| (0.200) | (0.169) | (0.161) | (0.177) | (0.157) | (0.149) | |
| Low Carbon Exp. pcTER | 5.945∗∗∗ | 13.126 | 34.828∗∗∗ | 7.612∗∗∗ | 23.200∗∗ | 29.144∗∗∗ |
| (1.559) | (12.028) | (11.466) | (1.383) | (11.155) | (10.604) | |
| Low Carbon Exp. pcLDR | 0.792∗∗∗ | 0.768∗∗∗ | 0.588∗∗∗ | 0.394∗∗∗ | 0.439∗∗∗ | 0.387∗∗∗ |
| (0.129) | (0.108) | (0.103) | (0.114) | (0.100) | (0.095) | |
| Low Carbon Exp. pcLDR | 6.004∗∗∗ | 36.630∗∗∗ | 29.030∗∗∗ | 3.151∗∗∗ | 18.037∗∗∗ | 13.218∗∗∗ |
| (0.736) | (5.048) | (4.862) | (0.653) | (4.682) | (4.497) | |
| GDP | 0.246∗∗∗ | 0.321∗∗∗ | 0.615∗∗∗ | 0.510∗∗∗ | ||
| (0.017) | (0.019) | (0.015) | (0.017) | |||
| EPS | 0.010∗∗ | 0.013∗∗∗ | 0.005 | 0.010∗∗ | ||
| (0.005) | (0.005) | (0.004) | (0.004) | |||
| Political Ideology | 0.017∗∗∗ | 0.001 | 0.002 | 0.007∗∗ | ||
| (0.004) | (0.004) | (0.003) | (0.003) | |||
| GVAA | 0.045∗∗∗ | 0.072∗∗∗ | ||||
| (0.013) | (0.012) | |||||
| GVAB-E | 0.222∗∗∗ | 0.084∗∗∗ | ||||
| (0.015) | (0.014) | |||||
| GVAF | 0.023 | 0.095∗∗∗ | ||||
| (0.014) | (0.013) | |||||
| GVAG-J | 0.0004 | 0.110∗∗∗ | ||||
| (0.030) | (0.027) | |||||
| GVAK-N | 0.311∗∗∗ | 0.498∗∗∗ | ||||
| (0.027) | (0.025) | |||||
| GVAO-U | 0.345∗∗∗ | 0.334∗∗∗ | ||||
| (0.033) | (0.030) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.510 | 0.614 | 0.653 | 0.787 | 0.796 | 0.817 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
| Dependent variable: | ||||||
| GHGEpc | GHGEi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.049 | 0.020 | 0.013 | 0.031 | 0.024 | 0.030 |
| (0.067) | (0.067) | (0.067) | (0.067) | (0.067) | (0.068) | |
| Low Carbon Exp. pc | 0.378 | 6.016∗ | 5.025 | 0.306 | 5.767∗ | 4.607 |
| (0.456) | (3.122) | (3.120) | (0.453) | (3.120) | (3.160) | |
| Low Carbon Exp. pcTER | 0.163 | 0.376∗∗∗ | 0.426∗∗∗ | 0.162 | 0.385∗∗∗ | 0.453∗∗∗ |
| (0.127) | (0.131) | (0.131) | (0.126) | (0.130) | (0.132) | |
| Low Carbon Exp. pcTER | 0.076 | 25.584∗∗∗ | 30.159∗∗∗ | 0.072 | 25.857∗∗∗ | 23.884∗∗ |
| (0.984) | (9.360) | (9.338) | (0.978) | (9.354) | (9.457) | |
| Low Carbon Exp. pcLDR | 0.028 | 0.001 | 0.009 | 0.044 | 0.025 | 0.013 |
| (0.082) | (0.085) | (0.085) | (0.082) | (0.085) | (0.086) | |
| Low Carbon Exp. pcLDR | 0.225 | 5.407 | 5.581 | 0.230 | 5.182 | 2.318 |
| (0.471) | (3.963) | (3.987) | (0.468) | (3.961) | (4.038) | |
| GDP | 0.460∗∗∗ | 0.519∗∗∗ | 0.535∗∗∗ | 0.464∗∗∗ | ||
| (0.028) | (0.035) | (0.028) | (0.035) | |||
| EPS | 0.026∗∗∗ | 0.031∗∗∗ | 0.018∗∗∗ | 0.018∗∗∗ | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| Political Ideology | 0.003 | 0.003 | 0.0004 | 0.0004 | ||
| (0.003) | (0.003) | (0.003) | (0.003) | |||
| GVAA | 0.010 | 0.007 | ||||
| (0.008) | (0.008) | |||||
| GVAB-E | 0.115∗∗∗ | 0.078∗∗∗ | ||||
| (0.016) | (0.016) | |||||
| GVAF | 0.086∗∗∗ | 0.030∗∗ | ||||
| (0.012) | (0.013) | |||||
| GVAG-J | 0.054∗∗ | 0.203∗∗∗ | ||||
| (0.025) | (0.025) | |||||
| GVAK-N | 0.076∗∗ | 0.134∗∗∗ | ||||
| (0.032) | (0.033) | |||||
| GVAO-U | 0.176∗∗∗ | 0.013 | ||||
| (0.037) | (0.037) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.191 | 0.235 | 0.241 | 0.103 | 0.078 | 0.061 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
Regression by type of polluting gas: trend components
Trend components
| Dependent variable: | ||||||
| CO2pc | CO2i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.170∗∗ | 0.161∗∗ | 0.002 | 0.172∗∗ | 0.193∗∗∗ | 0.094 |
| (0.078) | (0.066) | (0.062) | (0.073) | (0.062) | (0.058) | |
| Low Carbon Exp. pc | 0.678∗∗∗ | 0.833∗ | 0.234 | 0.926∗∗∗ | 1.337∗∗∗ | 0.617 |
| (0.255) | (0.481) | (0.453) | (0.239) | (0.456) | (0.427) | |
| Low Carbon Exp. pcTER | 0.003 | 0.268∗∗ | 0.013 | 0.084 | 0.097 | 0.110 |
| (0.149) | (0.134) | (0.126) | (0.140) | (0.127) | (0.118) | |
| Low Carbon Exp. pcTER | 1.718∗∗∗ | 0.049 | 2.866∗∗∗ | 2.102∗∗∗ | 0.847 | 2.222∗∗∗ |
| (0.501) | (0.947) | (0.894) | (0.470) | (0.899) | (0.842) | |
| Low Carbon Exp. pcLDR | 0.739∗∗∗ | 0.647∗∗∗ | 0.490∗∗∗ | 0.370∗∗∗ | 0.383∗∗∗ | 0.315∗∗∗ |
| (0.096) | (0.083) | (0.078) | (0.090) | (0.078) | (0.074) | |
| Low Carbon Exp. pcLDR | 2.603∗∗∗ | 3.441∗∗∗ | 2.423∗∗∗ | 1.514∗∗∗ | 2.267∗∗∗ | 1.621∗∗∗ |
| (0.255) | (0.494) | (0.466) | (0.239) | (0.469) | (0.439) | |
| GDP | 0.225∗∗∗ | 0.375∗∗∗ | 0.615∗∗∗ | 0.446∗∗∗ | ||
| (0.019) | (0.022) | (0.018) | (0.021) | |||
| EPS | 0.008 | 0.010∗∗ | 0.004 | 0.009∗ | ||
| (0.005) | (0.005) | (0.005) | (0.005) | |||
| Political Ideology | 0.022∗∗∗ | 0.002 | 0.008∗ | 0.006 | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| GVAA | 0.065∗∗∗ | 0.095∗∗∗ | ||||
| (0.014) | (0.014) | |||||
| GVAB-E | 0.244∗∗∗ | 0.055∗∗∗ | ||||
| (0.017) | (0.016) | |||||
| GVAF | 0.005 | 0.074∗∗∗ | ||||
| (0.016) | (0.015) | |||||
| GVAG-J | 0.032 | 0.143∗∗∗ | ||||
| (0.033) | (0.031) | |||||
| GVAK-N | 0.406∗∗∗ | 0.585∗∗∗ | ||||
| (0.031) | (0.029) | |||||
| GVAO-U | 0.430∗∗∗ | 0.435∗∗∗ | ||||
| (0.037) | (0.035) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.531 | 0.609 | 0.659 | 0.760 | 0.763 | 0.796 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
| Dependent variable: | ||||||
| CH4pc | CH4i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.134∗ | 0.100 | 0.103 | 0.136∗ | 0.132∗∗ | 0.199∗∗∗ |
| (0.079) | (0.064) | (0.064) | (0.075) | (0.065) | (0.066) | |
| Low Carbon Exp. pc | 0.378 | 0.077 | 0.114 | 0.130 | 0.580 | 0.965∗∗ |
| (0.258) | (0.471) | (0.468) | (0.247) | (0.473) | (0.483) | |
| Low Carbon Exp. pcTER | 0.037 | 0.033 | 0.044 | 0.044 | 0.138 | 0.080 |
| (0.151) | (0.131) | (0.130) | (0.145) | (0.132) | (0.134) | |
| Low Carbon Exp. pcTER | 2.751∗∗∗ | 2.460∗∗∗ | 3.132∗∗∗ | 3.133∗∗∗ | 3.258∗∗∗ | 2.488∗∗∗ |
| (0.507) | (0.928) | (0.923) | (0.487) | (0.933) | (0.952) | |
| Low Carbon Exp. pcLDR | 0.406∗∗∗ | 0.429∗∗∗ | 0.352∗∗∗ | 0.038 | 0.165∗∗ | 0.177∗∗ |
| (0.097) | (0.081) | (0.081) | (0.093) | (0.081) | (0.083) | |
| Low Carbon Exp. pcLDR | 0.404 | 1.339∗∗∗ | 0.836∗ | 0.683∗∗∗ | 0.163 | 0.034 |
| (0.258) | (0.484) | (0.481) | (0.248) | (0.486) | (0.496) | |
| GDP | 0.172∗∗∗ | 0.100∗∗∗ | 0.668∗∗∗ | 0.722∗∗∗ | ||
| (0.020) | (0.021) | (0.019) | (0.021) | |||
| EPS | 0.026∗∗∗ | 0.019∗∗∗ | 0.030∗∗∗ | 0.021∗∗∗ | ||
| (0.005) | (0.005) | (0.005) | (0.005) | |||
| Political Ideology | 0.003 | 0.003 | 0.017∗∗∗ | 0.011∗∗ | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| GVAA | 0.023 | 0.007 | ||||
| (0.015) | (0.015) | |||||
| GVAB-E | 0.108∗∗∗ | 0.191∗∗∗ | ||||
| (0.018) | (0.018) | |||||
| GVAF | 0.136∗∗∗ | 0.214∗∗∗ | ||||
| (0.016) | (0.017) | |||||
| GVAG-J | 0.112∗∗∗ | 0.001 | ||||
| (0.034) | (0.036) | |||||
| GVAK-N | 0.010 | 0.189∗∗∗ | ||||
| (0.032) | (0.033) | |||||
| GVAO-U | 0.170∗∗∗ | 0.175∗∗∗ | ||||
| (0.038) | (0.039) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.345 | 0.473 | 0.490 | 0.683 | 0.716 | 0.711 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
| Dependent variable: | ||||||
| N2Opc | N2Oi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.108 | 0.060 | 0.013 | 0.102∗ | 0.082 | 0.111∗∗ |
| (0.071) | (0.055) | (0.054) | (0.061) | (0.053) | (0.055) | |
| Low Carbon Exp. pc | 0.189 | 0.052 | 0.553 | 0.366∗ | 0.452 | 0.291 |
| (0.233) | (0.403) | (0.397) | (0.199) | (0.386) | (0.402) | |
| Low Carbon Exp. pcTER | 0.254∗ | 0.036 | 0.104 | 0.353∗∗∗ | 0.240∗∗ | 0.021 |
| (0.136) | (0.113) | (0.110) | (0.116) | (0.108) | (0.111) | |
| Low Carbon Exp. pcTER | 1.804∗∗∗ | 1.333∗ | 2.032∗∗∗ | 2.275∗∗∗ | 2.352∗∗∗ | 1.364∗ |
| (0.457) | (0.795) | (0.781) | (0.391) | (0.760) | (0.792) | |
| Low Carbon Exp. pcLDR | 0.693∗∗∗ | 0.344∗∗∗ | 0.290∗∗∗ | 0.331∗∗∗ | 0.085 | 0.116∗ |
| (0.087) | (0.069) | (0.068) | (0.075) | (0.066) | (0.069) | |
| Low Carbon Exp. pcLDR | 2.399∗∗∗ | 1.722∗∗∗ | 1.273∗∗∗ | 1.297∗∗∗ | 0.534 | 0.465 |
| (0.233) | (0.414) | (0.407) | (0.199) | (0.396) | (0.412) | |
| GDP | 0.103∗∗∗ | 0.135∗∗∗ | 0.759∗∗∗ | 0.714∗∗∗ | ||
| (0.018) | (0.019) | (0.015) | (0.018) | |||
| EPS | 0.006 | 0.010∗∗ | 0.010∗∗ | 0.012∗∗ | ||
| (0.004) | (0.004) | (0.004) | (0.005) | |||
| Political Ideology | 0.016∗∗∗ | 0.010∗∗∗ | 0.003 | 0.001 | ||
| (0.004) | (0.004) | (0.003) | (0.004) | |||
| GVAA | 0.091∗∗∗ | 0.062∗∗∗ | ||||
| (0.013) | (0.013) | |||||
| GVAB-E | 0.113∗∗∗ | 0.186∗∗∗ | ||||
| (0.015) | (0.015) | |||||
| GVAF | 0.040∗∗∗ | 0.039∗∗∗ | ||||
| (0.014) | (0.014) | |||||
| GVAG-J | 0.009 | 0.101∗∗∗ | ||||
| (0.029) | (0.030) | |||||
| GVAK-N | 0.188∗∗∗ | 0.369∗∗∗ | ||||
| (0.027) | (0.027) | |||||
| GVAO-U | 0.032 | 0.042 | ||||
| (0.036) | (0.036) | |||||
| Observations | 3,856 | 2,982 | 2,982 | 3,856 | 2,982 | 2,982 |
| Adjusted R2 | 0.109 | 0.159 | 0.203 | 0.594 | 0.619 | 0.595 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
Cycle components
| Dependent variable: | ||||||
| CO2pc | CO2i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.072 | 0.003 | 0.008 | 0.059 | 0.001 | 0.004 |
| (0.054) | (0.054) | (0.054) | (0.054) | (0.054) | (0.054) | |
| Low Carbon Exp. pc | 0.162 | 0.482 | 0.394 | 0.152 | 0.463 | 0.315 |
| (0.177) | (0.395) | (0.394) | (0.177) | (0.396) | (0.397) | |
| Low Carbon Exp. pcTER | 0.096 | 0.308∗∗∗ | 0.374∗∗∗ | 0.089 | 0.311∗∗∗ | 0.377∗∗∗ |
| (0.104) | (0.108) | (0.108) | (0.103) | (0.108) | (0.108) | |
| Low Carbon Exp. pcTER | 0.115 | 1.745∗∗ | 1.918∗∗ | 0.134 | 1.871∗∗ | 1.668∗∗ |
| (0.344) | (0.763) | (0.760) | (0.342) | (0.763) | (0.765) | |
| Low Carbon Exp. pcLDR | 0.062 | 0.052 | 0.063 | 0.069 | 0.067 | 0.067 |
| (0.067) | (0.068) | (0.068) | (0.067) | (0.068) | (0.068) | |
| Low Carbon Exp. pcLDR | 0.024 | 0.681∗ | 0.721∗ | 0.030 | 0.674∗ | 0.494 |
| (0.179) | (0.407) | (0.407) | (0.178) | (0.408) | (0.410) | |
| GDP | 0.567∗∗∗ | 0.660∗∗∗ | 0.428∗∗∗ | 0.323∗∗∗ | ||
| (0.034) | (0.042) | (0.034) | (0.042) | |||
| EPS | 0.036∗∗∗ | 0.041∗∗∗ | 0.028∗∗∗ | 0.028∗∗∗ | ||
| (0.004) | (0.005) | (0.004) | (0.005) | |||
| Political Ideology | 0.001 | 0.0004 | 0.003 | 0.002 | ||
| (0.003) | (0.003) | (0.003) | (0.003) | |||
| GVAA | 0.006 | 0.010 | ||||
| (0.010) | (0.010) | |||||
| GVAB-E | 0.125∗∗∗ | 0.069∗∗∗ | ||||
| (0.019) | (0.019) | |||||
| GVAF | 0.100∗∗∗ | 0.044∗∗∗ | ||||
| (0.015) | (0.015) | |||||
| GVAG-J | 0.078∗∗∗ | 0.179∗∗∗ | ||||
| (0.030) | (0.030) | |||||
| GVAK-N | 0.125∗∗∗ | 0.087∗∗ | ||||
| (0.039) | (0.039) | |||||
| GVAO-U | 0.291∗∗∗ | 0.103∗∗ | ||||
| (0.044) | (0.044) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.206 | 0.261 | 0.270 | 0.044 | 0.042 | 0.041 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
| Dependent variable: | ||||||
| CH4pc | CH4i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.198∗∗∗ | 0.174∗∗∗ | 0.184∗∗∗ | 0.185∗∗∗ | 0.173∗∗∗ | 0.172∗∗∗ |
| (0.047) | (0.048) | (0.048) | (0.047) | (0.047) | (0.048) | |
| Low Carbon Exp. pc | 0.075 | 0.188 | 0.168 | 0.085 | 0.207 | 0.247 |
| (0.152) | (0.348) | (0.348) | (0.152) | (0.346) | (0.353) | |
| Low Carbon Exp. pcTER | 0.152∗ | 0.237∗∗ | 0.256∗∗∗ | 0.146 | 0.240∗∗ | 0.259∗∗∗ |
| (0.089) | (0.095) | (0.095) | (0.089) | (0.094) | (0.096) | |
| Low Carbon Exp. pcTER | 0.699∗∗ | 1.353∗∗ | 1.455∗∗ | 0.718∗∗ | 1.479∗∗ | 1.206∗ |
| (0.295) | (0.671) | (0.670) | (0.294) | (0.667) | (0.681) | |
| Low Carbon Exp. pcLDR | 0.094 | 0.066 | 0.069 | 0.086 | 0.052 | 0.066 |
| (0.058) | (0.060) | (0.060) | (0.057) | (0.060) | (0.061) | |
| Low Carbon Exp. pcLDR | 0.199 | 0.198 | 0.124 | 0.194 | 0.190 | 0.105 |
| (0.153) | (0.359) | (0.359) | (0.153) | (0.356) | (0.364) | |
| GDP | 0.163∗∗∗ | 0.181∗∗∗ | 0.831∗∗∗ | 0.802∗∗∗ | ||
| (0.029) | (0.037) | (0.029) | (0.037) | |||
| EPS | 0.004 | 0.008∗∗ | 0.004 | 0.006 | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| Political Ideology | 0.002 | 0.001 | 0.002 | 0.001 | ||
| (0.003) | (0.003) | (0.003) | (0.003) | |||
| GVAA | 0.009 | 0.007 | ||||
| (0.009) | (0.009) | |||||
| GVAB-E | 0.048∗∗∗ | 0.146∗∗∗ | ||||
| (0.016) | (0.017) | |||||
| GVAF | 0.074∗∗∗ | 0.018 | ||||
| (0.013) | (0.013) | |||||
| GVAG-J | 0.002 | 0.260∗∗∗ | ||||
| (0.026) | (0.027) | |||||
| GVAK-N | 0.011 | 0.223∗∗∗ | ||||
| (0.034) | (0.035) | |||||
| GVAO-U | 0.045 | 0.233∗∗∗ | ||||
| (0.039) | (0.039) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.034 | 0.040 | 0.032 | 0.258 | 0.243 | 0.216 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
| Dependent variable: | ||||||
| N2Opc | N2Oi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.059 | 0.035 | 0.062 | 0.045 | 0.032 | 0.049 |
| (0.056) | (0.063) | (0.063) | (0.055) | (0.063) | (0.064) | |
| Low Carbon Exp. pc | 0.050 | 0.160 | 0.157 | 0.062 | 0.145 | 0.076 |
| (0.181) | (0.462) | (0.461) | (0.180) | (0.458) | (0.465) | |
| Low Carbon Exp. pcTER | 0.072 | 0.076 | 0.101 | 0.080 | 0.074 | 0.098 |
| (0.106) | (0.126) | (0.126) | (0.105) | (0.125) | (0.127) | |
| Low Carbon Exp. pcTER | 0.245 | 0.718 | 1.139 | 0.263 | 0.844 | 0.893 |
| (0.351) | (0.891) | (0.889) | (0.348) | (0.883) | (0.896) | |
| Low Carbon Exp. pcLDR | 0.191∗∗∗ | 0.214∗∗∗ | 0.193∗∗ | 0.184∗∗∗ | 0.200∗∗ | 0.190∗∗ |
| (0.069) | (0.080) | (0.080) | (0.068) | (0.079) | (0.080) | |
| Low Carbon Exp. pcLDR | 0.065 | 0.050 | 0.016 | 0.070 | 0.057 | 0.250 |
| (0.183) | (0.476) | (0.476) | (0.181) | (0.472) | (0.480) | |
| GDP | 0.140∗∗∗ | 0.157∗∗∗ | 0.856∗∗∗ | 0.826∗∗∗ | ||
| (0.035) | (0.049) | (0.035) | (0.049) | |||
| EPS | 0.0002 | 0.004 | 0.007 | 0.017∗∗∗ | ||
| (0.005) | (0.005) | (0.005) | (0.005) | |||
| Political Ideology | 0.009∗∗ | 0.007∗ | 0.005 | 0.005 | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| GVAA | 0.029∗∗ | 0.012 | ||||
| (0.012) | (0.012) | |||||
| GVAB-E | 0.039∗ | 0.156∗∗∗ | ||||
| (0.022) | (0.022) | |||||
| GVAF | 0.019 | 0.038∗∗ | ||||
| (0.017) | (0.018) | |||||
| GVAG-J | 0.114∗∗∗ | 0.147∗∗∗ | ||||
| (0.035) | (0.035) | |||||
| GVAK-N | 0.166∗∗∗ | 0.383∗∗∗ | ||||
| (0.046) | (0.046) | |||||
| GVAO-U | 0.031 | 0.154∗∗∗ | ||||
| (0.052) | (0.052) | |||||
| Observations | 3,856 | 2,982 | 2,982 | 3,856 | 2,982 | 2,982 |
| Adjusted R2 | 0.018 | 0.011 | 0.020 | 0.241 | 0.175 | 0.156 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||